OpenStudy (gabby1214):

I really don't understand this find the slope of the line that passes through the pair of points. (-41.72, 4.8) and (2.72,3.3) The slope is_?(integer or simplified fraction)

zepdrix (zepdrix):

\[\large\rm (x_1,y_1)=(-41.72,~4.8)\]\[\large\rm (x_2,y_2)=(2.72,~3.3)\] Calculate the difference(distance) between y1 and y2,\[\large\rm \Delta y=y_2-y_1\]\[\large\rm \Delta y=3.3-4.8\]

zepdrix (zepdrix):

Do the same for the x's, figure out how far apart they are,\[\large\rm \Delta x=x_2-x_1\]\[\large\rm \Delta x=2.72--41.72\]

zepdrix (zepdrix):

And then realize that slope is simply this rate of growth, it's the change in y, divided by the change in x. Slope measures how much we go up, divided by how much we go to the right.\[\large\rm m=\frac{\Delta y}{\Delta x}\]If the triangle is confusing, you can ignore that :)

zepdrix (zepdrix):

Gabbsssss :U